# Photodynamic melting of phase-reversed charge stripes and enhanced condensation

### v1

@@ Abstract
The non-linear optical coupling.—  That resonant targeting cannot couple to the target state does not impede higher-order, and thus nonlinear, optical excitation from taking place. This is particularly possible because intermediate equilibrium states can have a relatively large current matrix element, e.g., $|\langle 0|\hat{J}_{x}|6\rangle|\simeq 0.66$ and $|\langle 6|\hat{J}_{x}|2\rangle|\simeq 0.58$. Indeed, in second-order time-dependent perturbation theory, the time-dependent expansion coefficients in the interaction picture entering in $|\psi_{\rm I}(t)\rangle=\sum_{m}c_{m}(t)|m\rangle$, for initial state $|0\rangle$, final state $|2\rangle$, intermediate state $|m_{1}\rangle$, at infinite time, are (see Ref. [77])
@@
where $\hat{V}_{m,m^{\prime}}^{(n)}$ are the matrix elements of the $n$-th Fourier component of the perturbation in the unperturbed eigenstates $|m\rangle$ of $\hat{\cal H}$, and the delta function enforces total energy conservation via the absorption or emission of $n_{1}+n_{2}$ photons. Now, considering the intermediate state $|m_{1}=6\rangle$ has a significant contribution, a close resonance occurs for $E_{6}\approx E_{0}+n_{1}\Omega$, when the denominator in Eq. (3) is small. Meanwhile, the energy conservation condition yields $E_{2}=E_{0}+(n_{1}+n_{2})\Omega$. Figure 5(a) shows that at pump frequencies $\Omega=\omega_{62}\equiv E_{6}-E_{2}$, a large overlap with the target state $|2\rangle$ is obtained at long times. Using the two previous energy conditions, $\omega_{62}=-n_{2}\Omega$, indicating that $n_{2}=-1$ (a photon emission in between the intermediate and final states), explains the observations. As a result, the resonance condition leads to $E_{6}\approx E_{0}+n_{1}\omega_{62}$, therefore, $\omega_{60}/\omega_{62}\approx n_{1}$. The equilibrium eigenspectrum yields $\omega_{60}/\omega_{62}\simeq 2.75$, indicating that the resonant excitation process from the ground-state to the intermediate one likely involves the absorption of three quanta of energy $\Omega=\omega_{62}$.
@@
Lastly, to confirm this perturbation picture, the perturbation matrix elements explain an enlarged overlap at long times with other states as well, but at first-order: a considerable overlap with state $|6\rangle$ [Fig. 5(c)] is obtained at the resonance $\omega_{60}/\Omega_{\rm pump}=n_{1}=4.69/1.57\simeq 3$, via a three quanta absorption. In turn, the most prominent matrix element of $\hat{J}_{x}$ in the low-lying spectrum, $|\langle 3|\hat{J}_{x}|0\rangle|\simeq 1.93$, can be dynamically accessed in first-order with pump frequencies $\Omega_{\rm pump}\simeq\omega_{30}$ — Figure 5(b) precisely confirms this scenario.
Summary and outlook.—  We showed that fine-tuned photoirradiation, effectively promoting a non-linear optical excitation towards a target state, forms a pathway for dynamically tailoring a system to exhibit desired properties. This was exemplified in the case of the competing interplay between pairing and density-wave with phase reversal in the presence of stripes; however, the approach is likely applicable in other scenarios. Key is the identification of the equilibrium properties of the low-lying eigenspectrum, and the design of a perturbation that, in sufficiently low order, can access them. Notably, that generically melting stripes leads to transient superconductivity is a result under intense debate in experiments [78]. Our results affirmatively point to this connection in a specific model. Confirming this for the fermionic Hubbard and $t$-$J$ models relevant to the cuprates, and larger system sizes, requires numerics currently beyond the capabilities of exact diagonalization and Krylov methods we employ. This is currently only possible with other techniques that approximate the dynamics, such as the time-dependent variational principle in the language of matrix product states [79, 80, 81, 82] — we leave this for future investigation.
@@ Acknowledgements.
R.M. acknowledges support from the TcSUH Welch Professorship Award. R.T.S. is supported by the grant DOE DE-SC0014671 funded by the U.S. Department of Energy, Office of Science. Numerical simulations were performed in the Tianhe-2JK at the Beijing Computational Science Research Center and with resources provided by the Research Computing Data Core at the University of Houston. This work also used TAMU ACES at Texas A&M HPRC through allocation PHY240046 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296. The data that support the findings of this article are openly available [83].
Supplementary Materials
Photodynamic melting of phase-reversed charge stripes and enhanced condensation
Supplementary Materials
Photodynamic melting of phase-reversed charge stripes and enhanced condensation
These supplementary materials present additional results, including a detailed perturbative analysis of the photoirradiation, which justifies the investigation of matrix elements of current and kinetic energy operators. Analysis of such matrix elements also corroborates the choice of $x$-polarization, i.e., polarization perpendicular to the stripe directions. Additional investigations include the time evolution of the density during photodynamics and the usage of a pump resonant with other intermediate states.
1 Time-dependent Perturbation Theory for the photoirradiation
The main text explains the enhancement of zero-momentum occupancy and the melting of the charge $\pi$-phase shift by targeting eigenstates of the low-lying equilibrium energy spectrum. The manner by which this targeting is achieved is through the tuning of the pulse properties, in particular its frequency, such that at long times, the weight of the target state in the time-evolving wavefunction $|\psi(t)\rangle$ is maximized. We now explain how this arises as a consequence of time-dependent perturbation theory applied to the pulsed driving — Ref. [90] describes an equivalent analysis in another lattice model.
The effect of the pulse is to modify the equilibrium Hamiltonian $\hat{\cal H}$ [Eq. (1) of the main text] with a time-dependent perturbation $\hat{V}(t)$,
$\hat{H}(t)=\hat{\cal H}+\hat{V}(t)\ .$ (S1)
@@ 1 Time-dependent Perturbation Theory for the photoirradiation
$i\frac{d}{dt}|\psi_{\rm I}(t)\rangle=\hat{V}_{\rm I}(t)|\psi_{\rm I}(t)\rangle,$ (S3)
@@
$=|0\rangle+|\psi_{\rm I}^{(1)}(t)\rangle+|\psi_{\rm I}^{(2)}(t)\rangle+\ldots\ ,$ (S4)
@@
$|\psi_{\rm I}^{(1)}(t)\rangle=-i\int_{-\infty}^{t}dt_{1}\,\hat{V}_{\rm I}(t_{1})|0\rangle\ ,$ (S5)
@@
$|\psi_{\rm I}^{(2)}(t)\rangle=(-i)^{2}\int_{-\infty}^{t}dt_{2}\int_{-\infty}^{t_{2}}dt_{1}\,\hat{V}_{\rm I}(t_{2})\hat{V}_{\rm I}(t_{1})|0\rangle\ .$ (S6)
@@
$c_{m}^{(1)}(t)=-i\int_{-\infty}^{t}dt_{1}\,e^{iE_{m}t_{1}}\langle m|\hat{V}(t_{1})|0\rangle e^{-iE_{0}t_{1}}\ .$ (S7)
@@
$\times e^{iE_{m^{\prime}}t_{1}}\langle m^{\prime}|\hat{V}(t_{1})|0\rangle e^{-iE_{0}t_{1}}\ ,$ (S8)
@@
$\hat{V}(t)=\sum_{n=-\infty}^{\infty}\hat{V}_{n}e^{-in\Omega t},$ (S9)
then, in the infinite-time limit, the first-order transition amplitude [Eq.(S7)], after the time integration  [^7] Using the identity $\int_{-\infty}^{\infty}dte^{i\omega t}=2\pi\delta(\omega)$, becomes
$c_{m}^{(1)}(\infty)=-2\pi i\sum_{n}\langle m|\hat{V}_{n}|0\rangle\,\delta(E_{m}-E_{0}-n\Omega)\ ,$ (S10)
@@
$\times\delta(E_{m}-E_{0}-(n_{1}+n_{2})\Omega)\ ,$ (S11)
@@
Now that we have established the perturbative coefficients, we are in a position to describe their complete functional form. First, we recall the structure of $\hat{V}(t)$ in Eq. (S1), which arises from the hopping modification of the kinetic part of the Hamiltonian $\hat{H}(t)=\hat{K}(t)+\hat{V}(t)$]:
@@
$\quad\quad\quad\quad+i\sin A_{{\bf ij}}(t)\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{\bf j}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\Big{]}\ ,$ (S12)
where $A_{{\bf ij}}(t)={\bf A}(t)\cdot({\bf i}-{\bf j})$. As a result, the time-dependent perturbation in Eq. (S1) yields,
@@
$+i\sin A_{{\bf ij}}(t)$ $\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{\bf j}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\Big{]}\ ,$ (S13)
@@
$\hat{K}_{x}(t)\equiv-t_{h}\left(\cos A(t)-1\right)\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}+\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$ (S14)
@@
$\hat{J}_{x}(t)\equiv-it_{h}\sin A(t)\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\ .$ (S15)
Now there are two choices to achieve analytical expressions for the transition amplitudes: The first is to assume the perturbation is perfectly time-periodic, using the discrete Fourier transform [Eq. (S9)] to fully describe the Floquet-like driving of the system, i.e., by taking the pulse width $t_{d}\to\infty$, i.e., $A(t)\simeq A_{0}\cos(\Omega t)$ and the second is to taking into account the Gaussian envelope that modulates the sinusoidal pulse function. In the former fully periodic case, the discrete Fourier coefficients in Eq. S9, read
$\hat{V}_{n}=\frac{1}{T}\int_{0}^{T}dt\ \hat{V}(t)e^{in\Omega t},\quad\text{with,}\quad T=2\pi/\Omega.$ (S16)
@@
$=\sum_{m=-\infty}^{\infty}a_{m}e^{-im\Omega t}\ ,$ (S17)
@@
$=\sum_{m=-\infty}^{\infty}b_{m}e^{-im\Omega t}\ ,$ (S18)
where $\mathcal{J}_{n}(x)$ are the Bessel functions of the first kind of integer order $n$. After expanding $\cos(2n\Omega t)$ and $\sin((2n+1)\Omega t)$ to their exponential form, it leads to the identification $a_{0}={\cal J}_{0}(A_{0})-1$, $a_{\pm 2n}={\cal J}_{2n}(A_{0})$, $a_{\pm(2n+1)}=0$ and $b_{\pm(2n+1)}=\mp i{\cal J}_{2n+1}(A_{0})$, $b_{2n}=0$. This fully defines thus the terms entering in the transition amplitudes in Eqs. (S10) and (S12) as,
@@
$-it_{h}b_{n}\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\ .$ (S19)
In the second case, where $A(t)=A_{0}e^{-t^{2}/2t_{d}^{2}}\cos(\Omega t)$, instead of using the Fourier expansion  [^8] Which for this case should be instead a continuous Fourier transform unlike the discrete case of Eq. (S9)., insight can be gained by taking the limit of small pulse amplitudes, $A_{0}\ll 1$, such that $\cos A(t)\approx 1-\frac{1}{2}A^{2}(t)$ and $\sin A(t)\approx A(t)$, leading to
@@
$\equiv A(t)\hat{J}_{x}^{(2)}\ .$ (S21)
As a result, the first-order transition amplitude [Eq. (S7) for $t\to\infty$] becomes:
@@
$\quad+A_{0}^{2}\frac{\sqrt{\pi}}{2}\left[e^{-\frac{t_{d}^{2}}{4}\omega_{m0}^{2}}+\frac{1}{2}\left(e^{-\frac{t_{d}^{2}}{4}(\omega_{m0}-2\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{4}(\omega_{m0}+2\Omega)^{2}}\right)\right]\langle m|\hat{K}_{x}^{(2)}|0\rangle\Bigg{\}}\ ,$ (S22)
@@
$\quad+\left(\langle m|\hat{J}_{x}^{(2)}|m^{\prime}\rangle\langle m^{\prime}|\hat{K}_{x}^{(2)}|0\rangle+\text{H.c.}\right)\,\mathcal{I}_{JK}^{(m^{\prime})}\Big{]}\ ,$ (S23)
@@
$\quad+e^{-\frac{t_{d}^{2}}{2}\left[(\omega_{m^{\prime}0}+\Omega)^{2}+(\omega_{mm^{\prime}}-\Omega)^{2}\right]}\left.+e^{-\frac{t_{d}^{2}}{2}\left[(\omega_{m^{\prime}0}-\Omega)^{2}+(\omega_{mm^{\prime}}-\Omega)^{2}\right]}\right]\ ,$ (S24)
@@
$\quad\quad+e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}-2\Omega)^{2}+(\omega_{mm^{\prime}}+2\Omega)^{2}]}+e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}-2\Omega)^{2}+(\omega_{mm^{\prime}}-2\Omega)^{2}]}\biggr{]}\ ,$ (S25)
@@
$\quad+\frac{\pi t_{d}}{4}\left[e^{-\frac{t_{d}^{2}}{4}(\omega_{m^{\prime}0}-2\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{4}(\omega_{m^{\prime}0}+2\Omega)^{2}}\right]\left[e^{-\frac{t_{d}^{2}}{2}(\omega_{mm^{\prime}}-\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{2}(\omega_{mm^{\prime}}+\Omega)^{2}}\right]\ ,$ (S26)
@@
2 Relevant Matrix elements

### v2

@@ 6 The non-linear optical coupling
That resonant targeting cannot couple to the target state does not impede higher-order, and thus nonlinear, optical excitation from taking place. This is particularly possible because intermediate equilibrium states can have a relatively large current matrix element, e.g., $|\langle 0|\hat{J}_{x}|6\rangle|\simeq 0.66$ and $|\langle 6|\hat{J}_{x}|2\rangle|\simeq 0.58$. Indeed, in second-order time-dependent perturbation theory, the time-dependent expansion coefficients in the interaction picture entering in $|\psi_{\rm I}(t)\rangle=\sum_{m}c_{m}(t)|m\rangle$, for initial state $|0\rangle$, final state $|2\rangle$, intermediate state $|m_{1}\rangle$, at infinite time, are (see Appendix A)
@@
where $\hat{V}_{m,m^{\prime}}^{(n)}$ are the matrix elements of the $n$-th Fourier component of the perturbation in the unperturbed eigenstates $|m\rangle$ of $\hat{\cal H}$, and the delta function enforces total energy conservation via the absorption or emission of $n_{1}+n_{2}$ photons. Now, considering the intermediate state $|m_{1}=6\rangle$ has a significant contribution, a close resonance occurs for $E_{6}\approx E_{0}+n_{1}\Omega$, when the denominator in Eq. (3) is small. Meanwhile, the energy conservation condition yields $E_{2}=E_{0}+(n_{1}+n_{2})\Omega$. Figure 5(a) shows that at pump frequencies $\Omega=\omega_{62}\equiv E_{6}-E_{2}$, a large overlap with the target state $|2\rangle$ is obtained at long times. Using the two previous energy conditions, $\omega_{62}=-n_{2}\Omega$, indicating that $n_{2}=-1$ (a photon emission in between the intermediate and final states), explains the observations. As a result, the resonance condition leads to $E_{6}\approx E_{0}+n_{1}\omega_{62}$, therefore, $\omega_{60}/\omega_{62}\approx n_{1}$. The equilibrium eigenspectrum yields $\omega_{60}/\omega_{62}\simeq 2.75$, indicating that the resonant excitation process from the ground-state to the intermediate one likely involves the absorption of three quanta of energy $\Omega=\omega_{62}$. Resonance conditions with other intermediate states are explored in Appendix E.
@@
Lastly, to confirm this perturbation picture, the perturbation matrix elements explain an enlarged overlap at long times with other states as well, but at first-order: a considerable overlap with state $|6\rangle$ [Fig. 5(c)] is obtained at the resonance $\omega_{60}/\Omega_{\rm pump}=n_{1}=4.69/1.57\simeq 3$, via a three quanta absorption. In turn, the most prominent matrix element of $\hat{J}_{x}$ in the low-lying spectrum, $|\langle 3|\hat{J}_{x}|0\rangle|\simeq 1.93$, can be dynamically accessed in first-order with pump frequencies $\Omega_{\rm pump}\simeq\omega_{30}$ — Figure 5(b) precisely confirms this scenario. Analysis in terms of the symmetries of the operators involved in this perturbative picture further corroborates these results, and, together with the corresponding parity of the low-lying states targeted in the pump process, is detailed in Appendix H.
7 Summary and outlook
We showed that fine-tuned photoirradiation, effectively promoting a non-linear optical excitation towards a target state, forms a pathway for dynamically tailoring a system to exhibit desired properties. This was exemplified in the case of the competing interplay between pairing and density-wave with phase reversal in the presence of stripes; however, the approach is likely applicable in other scenarios. Key is the identification of the equilibrium properties of the low-lying eigenspectrum, and the design of a perturbation that, in sufficiently low order, can access them. Notably, that generically melting stripes leads to transient superconductivity is a result under intense debate in experiments [76]. Our results affirmatively point to this connection in a specific model. Confirming this for the fermionic Hubbard and $t$-$J$ models relevant to the cuprates, and larger system sizes, requires numerics currently beyond the capabilities of exact diagonalization and Krylov methods we employ. This is currently only possible with other techniques that approximate the dynamics, such as the time-dependent variational principle in the language of matrix product states [77, 78, 79, 80] — we leave this for future investigation.
@@ Acknowledgements.
R.M. acknowledges support from the TcSUH Welch Professorship Award. R.T.S. is supported by the grant DOE DE-SC0014671 funded by the U.S. Department of Energy, Office of Science. Numerical simulations were performed with resources provided by the Research Computing Data Core at the University of Houston. This work also used TAMU ACES at Texas A&M HPRC through allocation PHY240046 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296. The data that support the findings of this article are openly available [81].
Appendix A Time-dependent Perturbation Theory for the photoirradiation
The main text explains how to enhance zero-momentum occupancy and melt the charge $\pi$-phase shift by targeting eigenstates of the low-lying equilibrium energy spectrum. The manner by which this targeting is achieved is through the tuning of the pulse properties, in particular its frequency, such that at long times, the weight of the target state in the time-evolving wavefunction $|\psi(t)\rangle$ is maximized. We now explain how this arises as a consequence of time-dependent perturbation theory applied to the pulsed driving — Ref. [82] describes an equivalent analysis in another lattice model.
The effect of the pulse is to modify the equilibrium Hamiltonian $\hat{\cal H}$ [Eq. (1)] with a time-dependent perturbation $\hat{V}(t)$,
$\hat{H}(t)=\hat{\cal H}+\hat{V}(t)\ .$ (4)
@@ Appendix A Time-dependent Perturbation Theory for the photoirradiation
$i\frac{d}{dt}|\psi_{\rm I}(t)\rangle=\hat{V}_{\rm I}(t)|\psi_{\rm I}(t)\rangle,$ (6)
@@
$=|0\rangle+|\psi_{\rm I}^{(1)}(t)\rangle+|\psi_{\rm I}^{(2)}(t)\rangle+\ldots\ ,$ (7)
@@
$|\psi_{\rm I}^{(1)}(t)\rangle=-i\int_{-\infty}^{t}dt_{1}\,\hat{V}_{\rm I}(t_{1})|0\rangle\ ,$ (8)
@@
$|\psi_{\rm I}^{(2)}(t)\rangle=(-i)^{2}\int_{-\infty}^{t}dt_{2}\int_{-\infty}^{t_{2}}dt_{1}\,\hat{V}_{\rm I}(t_{2})\hat{V}_{\rm I}(t_{1})|0\rangle\ .$ (9)
@@
$c_{m}^{(1)}(t)=-i\int_{-\infty}^{t}dt_{1}\,e^{iE_{m}t_{1}}\langle m|\hat{V}(t_{1})|0\rangle e^{-iE_{0}t_{1}}\ .$ (10)
@@
$\times e^{iE_{m^{\prime}}t_{1}}\langle m^{\prime}|\hat{V}(t_{1})|0\rangle e^{-iE_{0}t_{1}}\ ,$ (11)
@@
$\hat{V}(t)=\sum_{n=-\infty}^{\infty}\hat{V}_{n}e^{-in\Omega t},$ (12)
then, in the infinite-time limit, the first-order transition amplitude [Eq.(10)], after the time integration  [^6] Using the identity $\int_{-\infty}^{\infty}dte^{i\omega t}=2\pi\delta(\omega)$, becomes
$c_{m}^{(1)}(\infty)=-2\pi i\sum_{n}\langle m|\hat{V}_{n}|0\rangle\,\delta(E_{m}-E_{0}-n\Omega)\ ,$ (13)
@@
$\times\delta(E_{m}-E_{0}-(n_{1}+n_{2})\Omega)\ ,$ (14)
@@
Now that we have established the perturbative coefficients, we are in a position to describe their complete functional form. First, we recall the structure of $\hat{V}(t)$ in Eq. (4), which arises from the hopping modification of the kinetic part of the Hamiltonian $\hat{H}(t)=\hat{K}(t)+\hat{V}(t)$]:
@@
$\quad\quad\quad\quad+i\sin A_{{\bf ij}}(t)\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{\bf j}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\Big]\ ,$ (15)
where $A_{{\bf ij}}(t)={\bf A}(t)\cdot({\bf i}-{\bf j})$. As a result, the time-dependent perturbation in Eq. (4) yields,
@@
$+i\sin A_{{\bf ij}}(t)$ $\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{\bf j}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\Big]\ ,$ (16)
@@
$\hat{K}_{x}(t)\equiv-t_{h}\left(\cos A(t)-1\right)\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}+\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$ (17)
@@
$\hat{J}_{x}(t)\equiv-it_{h}\sin A(t)\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\ .$ (18)
Now there are two choices to achieve analytical expressions for the transition amplitudes: The first is to assume the perturbation is perfectly time-periodic, using the discrete Fourier transform [Eq. (12)] to fully describe the Floquet-like driving of the system, i.e., by taking the pulse width $t_{d}\to\infty$, i.e., $A(t)\simeq A_{0}\cos(\Omega t)$ and the second is to taking into account the Gaussian envelope that modulates the sinusoidal pulse function. In the former fully periodic case, the discrete Fourier coefficients in Eq. 12, read
$\hat{V}_{n}=\frac{1}{T}\int_{0}^{T}dt\ \hat{V}(t)e^{in\Omega t},\quad\text{with,}\quad T=2\pi/\Omega.$ (19)
@@
$=\sum_{m=-\infty}^{\infty}a_{m}e^{-im\Omega t}\ ,$ (20)
@@
$=\sum_{m=-\infty}^{\infty}b_{m}e^{-im\Omega t}\ ,$ (21)
where $\mathcal{J}_{n}(x)$ are the Bessel functions of the first kind of integer order $n$. After expanding $\cos(2n\Omega t)$ and $\sin((2n+1)\Omega t)$ to their exponential form, it leads to the identification $a_{0}={\cal J}_{0}(A_{0})-1$, $a_{\pm 2n}={\cal J}_{2n}(A_{0})$, $a_{\pm(2n+1)}=0$ and $b_{\pm(2n+1)}=\mp i{\cal J}_{2n+1}(A_{0})$, $b_{2n}=0$. This fully defines thus the terms entering in the transition amplitudes in Eqs. (13) and (15) as,
@@
$-it_{h}b_{n}\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\ .$ (22)
In the second case, where $A(t)=A_{0}e^{-t^{2}/2t_{d}^{2}}\cos(\Omega t)$, instead of using the Fourier expansion  [^7] Which for this case should be instead a continuous Fourier transform unlike the discrete case of Eq. (12)., insight can be gained by taking the limit of small pulse amplitudes, $A_{0}\ll 1$, such that $\cos A(t)\approx 1-\frac{1}{2}A^{2}(t)$ and $\sin A(t)\approx A(t)$, leading to
@@
$\equiv A(t)\hat{J}_{x}^{(2)}\ .$ (24)
As a result, the first-order transition amplitude [Eq. (10) for $t\to\infty$] becomes:
@@
$\quad+A_{0}^{2}\frac{\sqrt{\pi}}{2}\left[e^{-\frac{t_{d}^{2}}{4}\omega_{m0}^{2}}+\frac{1}{2}\left(e^{-\frac{t_{d}^{2}}{4}(\omega_{m0}-2\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{4}(\omega_{m0}+2\Omega)^{2}}\right)\right]\langle m|\hat{K}_{x}^{(2)}|0\rangle\Bigg\}\ ,$ (25)
@@
$\quad+\left(\langle m|\hat{J}_{x}^{(2)}|m^{\prime}\rangle\langle m^{\prime}|\hat{K}_{x}^{(2)}|0\rangle+\text{H.c.}\right)\,\mathcal{I}_{JK}^{(m^{\prime})}\Big]\ ,$ (26)
@@
$\quad+e^{-\frac{t_{d}^{2}}{2}\left[(\omega_{m^{\prime}0}+\Omega)^{2}+(\omega_{mm^{\prime}}-\Omega)^{2}\right]}\left.+e^{-\frac{t_{d}^{2}}{2}\left[(\omega_{m^{\prime}0}-\Omega)^{2}+(\omega_{mm^{\prime}}-\Omega)^{2}\right]}\right]\ ,$ (27)
@@
$\quad\quad+e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}-2\Omega)^{2}+(\omega_{mm^{\prime}}+2\Omega)^{2}]}+e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}-2\Omega)^{2}+(\omega_{mm^{\prime}}-2\Omega)^{2}]}\biggr]\ ,$ (28)
@@
$\quad+\frac{\pi t_{d}}{4}\left[e^{-\frac{t_{d}^{2}}{4}(\omega_{m^{\prime}0}-2\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{4}(\omega_{m^{\prime}0}+2\Omega)^{2}}\right]\left[e^{-\frac{t_{d}^{2}}{2}(\omega_{mm^{\prime}}-\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{2}(\omega_{mm^{\prime}}+\Omega)^{2}}\right]\ ,$ (29)
@@
Appendix B Relevant Matrix elements